OpenStudy (anonymous):

Find cos Θ if sin Theta = a/2. Assume the terminal side of the angle falls in Quadrant 2. Show all work. Find sin Θ if tan Theta = d/4. Assume the terminal side of the angle falls in Quadrant 1. Show all work. For the second problem, I have worked out d/4 +1= 1/cos^2 theta But I am unable to figure out what to do from there on.

OpenStudy (anonymous):

\[\cos(\theta) = \sin(\theta+\frac\pi2) = \sin(\theta+90^\circ)\]

OpenStudy (loser66):

still need help?

OpenStudy (anonymous):

Yes please. I am still stuck on the second problem :)

OpenStudy (loser66):

you got the first one? show me your stuff, please . (the first one)

OpenStudy (anonymous):

Not exactly, I tried using what broken symmetry had put but I had gotten confused.  cos= -2/ 2 sqrt3 since cos is x/r

OpenStudy (loser66):

ok, I guide you step by step , for the first one, the triangle in quadrant 2 has the shape like this